I Love My Chocolate Company makes dark chocolate and light chocolate. Both products require cocoa and sugar. The following planning information has been made available:

Standard Amount per Case
Dark Chocolate Light Chocolate Standard Price per Pound
Cocoa 9 lb. 6 lb. $5.4
Sugar 7 lb. 11 lb. 0.6
Standard labor time 0.4 hr. 0.5 hr.
Dark Chocolate Light Chocolate
Planned production 5,600 cases 13,200 cases
Standard labor rate $16.5 per hr. $16.5 per hr.

I Love My Chocolate does not expect there to be any beginning or ending inventories of cocoa or sugar. At the end of the budget year, I Love My Chocolate Company had the following actual results:

Dark Chocolate Light Chocolate
Actual production (cases) 5,300 13,700
Actual Price per Pound Actual Pounds Purchased and Used
Cocoa $5.5 130,500
Sugar 0.55 183,100
Actual Labor Rate Actual Labor Hours Used
Dark chocolate $16.1 per hr. 1,930
Light chocolate 16.9 per hr. 7,020

Required:

Prepare the following variance analyses for both chocolates and total, based on the actual results and production levels at the end of the budget year:

  1. Direct materials price variance, direct materials quantity variance, and total variance.
  2. Direct labor rate variance, direct labor time variance, and total variance.

Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. If there is no variance, enter a zero.

a. Direct materials price variance $
Direct materials quantity variance $
Total direct materials cost variance $
b. Direct labor rate variance $
Direct labor time variance $
Total direct labor cost variance $

2. The variance analyses should be based on the   amounts at   volumes. The budget must flex with the volume changes. If the   volume is different from the planned volume, as it was in this case, then the budget used for performance evaluation should reflect the change in direct materials and direct labor that will be required for the   production. In this way, spending from volume changes can be separated from efficiency and price variances.
Answer #1
Cocoa
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($5.40/ib. - $5.50/ib ) × 130500 ib. = $13050 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((9 ibs. X 5300 Unit)+(6ibsX13700)-130500)X $5.40= $3240 U F
For Sugar
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($0.60 - $0.55 ) × 183100 = $9155F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((7 ibs. X 5300 Unit)+(11 ibsX13700)-183100)X $0.60= $2820 F
Total Material Price Variance= $13050 UF+$9155 F= $3895 UF
Total Material Quantity Variance= $3240 UF+ $2820 F=$420 UF
Total Material Cost Variance= $3895 UF+ $420 UF= $4315UF
Dark Choclate
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($16.5/hour - $16.10/hour) × 1930 Hours= $772 F
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard RAte
= (0.40 Hour × 5300 Unit - 1930) × $16.50 = $3135 F
Light Choclate
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($16.50/hour - $16.90/hour) × 7020 Hours= $2808 UF
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard overhead recovery rate
= (0.50 Hour × 13700 Unit - 7020) × $16.5 = $2805 U F
Total LAbour Rate Variance= $772 F+ $2808 UF= $2036 UF
Total Labour Efficiency Variance= $3135F+$2805 UF=$330 F
Total Labour Cost Variance = $2036 UF+ $330F=$1706 UF
The variance analyses should be based on the standard amounts at actual volumes. The budget must flex with the volume changes. If the actual volume is different from the planned volume, as it was in this case, then the budget used for performance evaluation should reflect the change in direct materials and direct labor that will be required for the actual production. In this way, spending from volume changes can be separated from efficiency and price variances.